Use technology to find the P-value for the hypothesis test described below. The claim is that for a smartphone carrier's data speeds at airports, the mean is μ=15.00 Mbps. The sample size is n=14 and the test statistic is t=1.176.

P-value=

(Round to three decimal places as needed.)
Answer #1

sal: Guiven, Ho! Mz 15 Mbps VS Heil & 15 Mbps n=14 it = 1.276 13 df= m-1 = This is two- tailed test. P-value is i- of = 13, t